OpenStudy (timaashorty):

Can anyone take me through a matrices question step by step? Please(:

OpenStudy (amistre64):

maybe ... whats the question?

OpenStudy (amistre64):

if tis QR decomp im a bit rusty

OpenStudy (timaashorty):

Find -5A + 4B

OpenStudy (timaashorty):

Do you know that?

OpenStudy (amistre64):

oh, something rather basic ... so whats giving you the issue?

OpenStudy (amistre64):

yeah .... its just multiplication and addition

OpenStudy (amistre64):

multiply all of the elements in A by -5, and all the elements in B by 4 then add them together by like 'components'

OpenStudy (timaashorty):

everything lol Idk where to start

OpenStudy (amistre64):

start with ... multiply all of the elements in A by -5 then move on to multiplying all the elements in B by 4 end by adding the results by like components

OpenStudy (timaashorty):

6*(-5) = -30 1(-5) = -5 -4(-5) = 20 -6(-5) = 30 7(-5) =-35 -7(-5) = 35 b. -5(4) = -20 -1(4) = -4 -3(4) = -12 -8(4) = -32 6(4) = 24 8(4) = 32

OpenStudy (timaashorty):

Like that?

OpenStudy (amistre64):

that looks good so far :)

OpenStudy (timaashorty):

I add it now?

OpenStudy (amistre64):

thats perfect, and a nice format really now add the first of A with the first of B then the second of A with the second of B and all the way to the end

OpenStudy (timaashorty):

Thanks, (: & bear with me im kinda slow lol

OpenStudy (amistre64):

6(-5) = -30 -5(4) = -20: -50 1(-5) = -5 -1(4) = -4: -9 -4(-5) = 20 -3(4) = -12: 8 -6(-5) = 30 -8(4) = -32: -2 7(-5) =-35 6(4) = 24: -11 -7(-5) = 35 8(4) = 32: 67

OpenStudy (amistre64):

youre doing fine :)

OpenStudy (timaashorty):

thanks for making it easier 50 + -9= 41 8 + -2 = 6 -11 + 67 = 56

OpenStudy (amistre64):

you only need to do the addition once ... the matrix solution will simply be \[-5A + 4B = \begin{vmatrix} -50 & -9\\ 8 & -2\\ -11& 67\\ \end{vmatrix}\]

OpenStudy (timaashorty):

Oh isee my bad, what do I do next?

OpenStudy (timaashorty):

Oh wait thats it right?

OpenStudy (amistre64):